40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 19 14:06:36 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_14-06-11.png
    Reply  Fri Apr 19 20:21:52 2019, gautam, Update, PSL, Innolight NPRO shutoff Screenshot_from_2019-04-19_20-27-04.png
       Reply  Wed Apr 24 16:06:44 2019, gautam, Update, PSL, Innolight NPRO shutoff PMCrefl.png
          Reply  Thu Apr 25 01:03:29 2019, gautam, Update, PSL, MC trans is ~1000 cts (~7%) lower than usual 
          Reply  Thu Apr 25 17:31:56 2019, gautam, Update, PSL, Innolight NPRO shutoff EdwinShutoff20190425.pngEdwinShutdown_zoomOut.png
             Reply  Mon Apr 29 16:25:22 2019, gautam, Update, PSL, PSL turned on again 
                Reply  Fri May 3 12:51:58 2019, gautam, Update, PSL, PSL turned on again 
                   Reply  Mon May 6 10:51:43 2019, gautam, Update, PSL, PSL turned off again NPROshutoff.png
                   Reply  Wed May 8 19:04:20 2019, rana, Update, PSL, PSL turned on again 
                      Reply  Thu May 9 19:50:04 2019, gautam, Update, PSL, PSL turned off again Screenshot_from_2019-05-09_19-49-29.png
                         Reply  Mon May 13 10:45:38 2019, gautam, Update, PSL, PSL turned ON again IMG_7431.JPG
                            Reply  Wed May 15 10:56:47 2019, gautam, Update, PSL, PSL turned ON again 
Message ID: 14597     Entry time: Wed May 8 19:04:20 2019     In reply to: 14593     Reply to this: 14599
Author: rana 
Type: Update 
Category: PSL 
Subject: PSL turned on again 
  1. Increased PSL HEPA Variac from 30 to 100% to get more airflow.
  2. All of the TEC setpoints seem cold to me, so I increased the laser crystal temperature to 30.6 C
  3. Adjusted the diode TEC setpoints individually to optimize the PMC REFL power (unlocked). DTEC A = 22.09 C, DTEC B = 21.04 C
  4. locked PMC at 1900 PT; let's see how long it lasts.

My hunch is that the TECs are working too hard and can't offload the heat onto the heat sinks. As the diode's degrade, more of the electrical power is converted to heat in the diodes rather than 808 nm photons. So hopefully the increased airflow will help.

I tried to increase the DTEC setpoints, but that seems to detune them too far from the laser absorption band, so that's not very efficient for us. IN any case, if we end up changin the laser temperature, we'll have to adjust the ALS lasers to match, and that will be annoying.

 

The office area was very cold and the HVAC air flow stronger than usual. I changed the setpoint on the thermostat near Steve's desk from 71 to 73F at 1830 today.

ELOG V3.1.3-